# Integrated modeling of boron powder injection for real-time plasma-facing component conditioning

### v2

@@ 7 Summary and conclusions
A new comprehensive modeling approach for real-time wall conditioning and in-situ boron powder coating is presented, utilizing three advanced models: EMC3-EIRENE, DIS, and WallDYN3D [18, 21, 25, 26, 27]. These models collectively address 3D plasma transport, powder particle migration and ablation, and mixed-material surface evolution. Ablation modeling with EMC3-EIRENE/DIS shows that boron particles of up to 150 $\mu$m evaporate at the separatrix, where they reach evaporation temperature. The transport analysis shows that initial B fluxes are higher on the inboard divertor target due to the directional frictional forces from the main ion flow in the evaporation region. This effect closely aligns with findings from EMC3-EIRENE simulations using a single atomic B source in the same plasma scenario previously [17], and is consistent with beryllium transport modeling for JET and ITER using ERO2.0 [36]. However, first-flight deposition based on EMC3-EIRENE/DIS alone neglects impurity recycling, resulting in significant asymmetries in the divertor B fluxes concentrated in a narrow, far SOL region. The powder particle mass was not found to affect the distribution of B fluxes between the inboard and outboard divertor targets for particle sizes 5-250 $\mu$m in the present plasma scenario.
The modeling was extended by taking additional mixed-material sources and sinks, i.e., erosion and redeposition, into account using the WallDYN3D code. The code was extended to include external impurity sources from powder injection. Quasi-equilibrium solutions obtained with WallDYN3D show a toroidally uniform distribution with B fluxes on the outboard divertor target exceeding those on the inboard target. The approach also replicates B surface concentrations on the outer divertor close to experimental findings. However, the model underpredicts the B fluxes necessary to achieve this balance, likely partially due to a limited plasma grid domain that mainly covers the divertor region and neglects potential additional C sources from the remaining wall. Therefore, extending the simulation grid for the whole wall will be considered crucial for future work.
@@ Acknowledgements
This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Fusion Energy Sciences, using the DIII-D National Fusion Facility, a DOE Office of Science user facility, under Awards DE-AC02-09CH11466 (PPPL), DE-FC02-04ER54698 (DIII-D), and DE-AC05-00OR22725 (ORNL). The United States Government retains a non-exclusive, paid-up, irrevocable, world-wide license to publish or reproduce the published form of this manuscript, or allow others to do so, for United States Government purposes. The data that support the findings of this study will be openly available in Princeton Data Commons at https://doi.org/10.34770/w32p-2w69.

### v3

@@ 7 Summary and conclusions
A new comprehensive modeling approach for real-time wall conditioning and in-situ boron powder coating is presented, utilizing three advanced models: EMC3-EIRENE, DIS, and WallDYN3D [18, 21, 25, 26, 27]. These models collectively address 3D plasma transport, powder particle migration and ablation, and mixed-material surface evolution. Ablation modeling with EMC3-EIRENE/DIS shows that boron particles of up to 150 $\mu$m evaporate at the separatrix. The EMC3 trace impurity fluid modeling shows that initial B fluxes are higher on the inboard divertor target due to the directional frictional forces from the main ion flow in the evaporation region. This effect aligns with findings from EMC3-EIRENE simulations using a single atomic B source in the same plasma scenario previously [17]. This trend is consistent with beryllium transport modeling for JET and ITER using ERO2.0 [36]. However, first-flight deposition based on EMC3-EIRENE/DIS alone neglects impurity recycling, resulting in significant asymmetries in the divertor B fluxes concentrated in a narrow, far SOL region. The powder particle mass was not found to affect the distribution of B fluxes between the inboard and outboard divertor targets for particle sizes 5-250 $\mu$m in the present plasma scenario.
Additional mixed-material sources and sinks, i.e., erosion, reflection, and redeposition, were taken into account using the WallDYN3D code. The code was extended to include external impurity sources from powder injection. The kinetic treatment of impurities in WallDYN3D initially shows a reduced imbalance in B fluxes between the inner and outer divertor targets compared to the EMC3 fluid model. B and C primarily recycle locally through erosion and redeposition, with the outer target showing a higher B influx due to greater erosion and reflection flux. This leads to near-zero net erosion of C and minor B accumulation on the outer target, supported by local recycling and the continuous B source, while the inner target shows net erosion of C and less local B redeposition. The surface composition is governed by the dynamic balance of erosion and redeposition rather than long-range transport. The approach also replicates B surface concentrations on the outer divertor close to experimental findings. However, the model underpredicts the B fluxes necessary to achieve this balance, likely partially due to a limited plasma grid domain that mainly covers the divertor region and neglects potential additional C sources from the remaining wall. Therefore, extending the simulation grid for the whole wall will be considered crucial for future work.
@@ Acknowledgements
This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Fusion Energy Sciences, using the DIII-D National Fusion Facility, a DOE Office of Science user facility, under Awards DE-AC02-09CH11466, DE-FC02-04ER54698, and DE-AC05-00OR22725. The United States Government retains a non-exclusive, paid-up, irrevocable, world-wide license to publish or reproduce the published form of this manuscript, or allow others to do so, for United States Government purposes. The data that support the findings of this study are openly available in Zenodo at https://doi.org/10.5281/zenodo.14176757.